40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 16 21:05:24 2015, rana, Update, General, some status 48.png39.png
    Reply  Mon May 18 09:50:00 2015, ericq, Update, General, some status 
       Reply  Mon May 18 11:59:07 2015, rana, Update, General, some status 
          Reply  Tue May 19 16:18:57 2015, ericq, Update, General, crons fixed 
    Reply  Mon May 18 16:28:18 2015, ericq, Update, General, some status 
       Reply  Mon May 18 18:03:12 2015, rana, Update, General, some status CPUtrend.png
          Reply  Tue May 19 00:19:23 2015, rana, Update, General, some status Screen_Shot_2015-05-19_at_12.17.39_AM.png
             Reply  Wed May 20 03:08:27 2015, rana, Update, General, some status 
    Reply  Mon May 18 17:42:14 2015, rana, ericQ, Update, General, some status 
       Reply  Tue May 19 18:55:12 2015, rana, ericQ, Update, General, some status 
          Reply  Wed May 20 11:41:59 2015, ericq, Update, General, some status 
    Reply  Mon Sep 21 00:51:36 2015, rana, Update, General, op340m, autoburt cron =? megatron 
       Reply  Mon Sep 21 11:40:30 2015, ericq, Update, General, Megatron maitenence 
          Reply  Tue Dec 1 17:26:14 2015, Koji, Update, General, Megatron maitenence 
             Reply  Tue Dec 1 20:20:16 2015, Koji, Update, General, Megatron maitenence 
Message ID: 11303     Entry time: Mon May 18 17:42:14 2015     In reply to: 11294     Reply to this: 11315
Author: rana, ericQ 
Type: Update 
Category: General 
Subject: some status 

Today at 5 PM we replaced the east N2 cylinder. The east pressure was 500 and the west cylinder pressure was 1000. Since Steve's elogs say that the consumption can be as high as 800 per day we wanted to be safe.

  1. We closed the black valve before the regulator and closed the valve on the cylinder.
  2. We unscrewed the brass fill line to the cylinder.
  3. We unchained the cylinder and put in the dolly (and attached the chains on there).
  4. We rolled in a fresh cylinder from outside using the red dolly (it should have chains).
  5. We put it in place, hooked up the chains, and screwed on the brass nozzle with the large adjustable wrench (need to put a non-adjustable here).
  6. Opened up the cylinder valve.
  7. Opened up the black valve.
  8. New east pressure reading is 2500 PSI. Regulated N2 pressure is 68 PSI.
Quote:

1) Checked the N2 pressures: the unregulated cylinder pressures are both around 1500 PSI. How long until they get to 1000?

 

ELOG V3.1.3-